How much money must you deposit into a savings account at the end of each year at 4% interest compounded annually in order to earn $9,778.08 interest during a 20-year period?
Answer #1

Let the amount deposited per annum be X.

Then future value (FV) of the deposit at the end of 20 years @ 4% p.a. is

X*[(1+4%)^20 -1]/4% =

Total amount deposited over 20 years = 20*X = 20X

Total interest earned = FV of deposit - total amount deposited

9,778.08 = 29.7781X -20X

X = 9,778.08/9.7781 = 1,000

Deposit has to be $1,000 per annum.